### v1

# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

Abstract
Multimodal contrastive learning models (e.g., CLIP) can learn high-quality representations from large-scale image-text datasets, yet they exhibit significant vulnerabilities to backdoor attacks, raising serious safety concerns. In this paper, we disclose that CLIP’s vulnerabilities primarily stem from its excessive encoding of class-irrelevant features, which can compromise the model’s visual feature resistivity to input perturbations, making it more susceptible to capturing the trigger patterns inserted by backdoor attacks. Inspired by this finding, we propose Repulsive Visual Prompt Tuning (RVPT), a novel defense approach that employs specially designed deep visual prompt tuning and feature-repelling loss to eliminate excessive class-irrelevant features while simultaneously optimizing cross-entropy loss to maintain clean accuracy. Unlike existing multimodal backdoor defense methods that typically require the availability of poisoned data or involve fine-tuning the entire model, RVPT leverages few-shot downstream clean samples and only tunes a small number of parameters. Empirical results demonstrate that RVPT tunes only 0.27% of the parameters relative to CLIP, yet it significantly outperforms state-of-the-art baselines, reducing the attack success rate from 67.53% to 2.76% against SoTA attacks and effectively generalizing its defensive capabilities across multiple datasets. Our anonymous code is available on https://anonymous.4open.science/r/RVPT.
1 Introduction
Figure 1: The empirical distribution of the cosine similarity between the penultimate-layer features/logits of the original images and those of its slightly, non-adversarially perturbed version. Additionally, we obtain the empirical distribution of the mean cosine similarity between the penultimate-layer features of each image and those of images in its class, referred to as intra-class image-image similarity. Four encoders are evaluated: (a) ViT pre-trained exclusively on ImageNet [40]. (b) ViT from openAI’s CLIP [41]. (c) ViT of the backdoored CLIP tuned with Repulsive Visual Prompt Tuning (RVPT) using 1 shot per class on ImageNet. (d) ViT of the backdoored CLIP tuned with RVPT using 16 shots per class on ImageNet. We also calculate each encoder’s Perturbation Resistivity (PR), which reflects the resistivity of the encoded features against input perturbations, and the Attack Success Rate (ASR) against the backdoored encoder, which indicates the encoder’s vulnerability to backdoor attack during inference. Specifically, CLIP in (c) and (d) is backdoored and attacked by BadCLIP [28].
Recently, Contrastive Language-Image Pretraining (CLIP) [41] has emerged as a powerful base model across multiple domains. Unlike traditional models that rely on uni-modal supervision, CLIP aligns visual and textual signals to learn rich representations from numerous image-text pairs in the pre-training stage. This unique training paradigm enables its impressive performance in open-vocabulary visual recognition [53] and robustness to distribution shift [12].
However, despite CLIP’s success in visual recognition, recent research [3, 4] has revealed its vulnerabilities against backdoor attacks in downstream tasks. The vulnerabilities can be roughly categorized into two aspects. First, concerns arise from its training data collection process: the data is noisy, uncurated, and sourced from the web, which facilitates adversarial poisoning of the image-text pairs. Second, CLIP demonstrates heightened susceptibility to a smaller number of poisoned training samples: recent studies [3] have empirically shown that poisoning CLIP may require injecting orders of magnitude fewer poisoned data into the training dataset than poisoning conventional supervised models. Once poisoned during pre-training, the backdoored CLIP will inevitably classify images with an adversarially designed trigger into the predefined target class.
CLIP’s vulnerabilities to backdoor attacks have prompted the development of various defense methods. However, most of these methods necessitate fine-tuning the entire model’s parameters [2] or require the availability of poisoned data [57, 19, 56], which could be resource-intensive or even unrealistic. Thus, to achieve an efficient, effective, and practical defense, we are dedicated to developing a defense method based on visual prompt tuning (VPT) [21], utilizing only a small portion of downstream clean samples to eliminate the embedded trigger feature directly. However, simply adopting VPT proves ineffective in experiments, as the visual encoder will not extract trigger features from clean samples, making it impossible for VPT to learn to remove them. Notably, we observe that CLIP’s visual features exhibit unusually high sensitivity to even non-adversarial input perturbations, a trait strongly associated with its significant vulnerability to backdoor attacks. Therefore, we can instead leverage VPT to lower this sensitivity and enhance backdoor robustness.
To quantify this sensitivity, we introduce the Perturbation Resistivity (PR) metric. PR is defined as the ratio of the expected cosine similarity between the penultimate-layer features of original and perturbed images and the expected cosine similarity between the logits of these images:
$\text{PR}(D,\boldsymbol{\delta})=\frac{\mathbb{E}_{(\boldsymbol{x},y)\sim D}[\cos(\phi(\boldsymbol{x}),\phi(\boldsymbol{x}+\boldsymbol{\delta}))]}{\mathbb{E}_{(\boldsymbol{x},y)\sim D}[\cos(g(\boldsymbol{x}),g(\boldsymbol{x}+\boldsymbol{\delta}))]}.$ (1)
Here, $\phi(\boldsymbol{x})$ represents the penultimate-layer features for the input $\boldsymbol{x}$, $g(\boldsymbol{x})$ denotes the logits for the input $\boldsymbol{x}$, $D$ represents the given dataset and $\cos(\cdot,\cdot)$ denotes cosine similarity. A lower PR value indicates a greater sensitivity of the encoded features to input perturbations, reflecting the model’s potential vulnerability against backdoor attacks due to increased susceptibility to capturing trigger patterns. In particular, we measured the ASRs for models with different PRs and observed a strong negative correlation between ASR and PR, which is presented in the supplemental material. Notably, unless otherwise noted, we abbreviate empirical PR as PR.
In order to investigate the PR values of different visual encoders, we randomly sampled 5,000 images from the ImageNet [9] validation set and introduced random noise [15] of 16$\times$16 pixels as the perturbation to these images. Results for more perturbations are provided in the supplemental material. We expect all the PR values to be close to 1 since the perturbation is neither adversarial nor semantic. However, the results depicted in Figure 1 (a) (b) show that the PR value of CLIP ViT (Vision Transformer [11]) is greatly lower than that of ViT trained on ImageNet, indicating CLIP’s high sensitivity to perturbations and vulnerability to backdoor attacks. To explain this phenomenon, we also analyze each model’s intra-class image-image similarity of the sampled images in Figure 1 and discover that CLIP demonstrates lower similarity compared with traditional supervised models, suggesting that CLIP encodes a broader range of class-irrelevant features, which do not correlate with any class in the specific downstream dataset. Therefore, CLIP’s tendency to excessively encode class-irrelevant features is supposed to be a key factor contributing to its lower PR, as it enables CLIP to capture and encode diverse and fine-grained patterns more easily. In summary, the unique training paradigm of CLIP results in its capture of a broad and potentially excessive range of class-irrelevant features, thereby heightening the sensitivity of its visual features to input perturbations and rendering the CLIP model vulnerable to backdoor attacks.
To address this challenge, we propose a novel approach termed Repulsive Visual Prompt Tuning (RVPT), to trim the class-irrelevant features (CIFs) in a few-shot manner. As shown in Figure 2, RVPT mainly utilizes deep visual prompt tuning to eliminate CIFs. Specifically, it adopts feature-repelling loss to repel the prompted features from the original features by minimizing their mean cosine similarity while optimizing cross-entropy loss to maintain clean accuracy. This mechanism ensures that CIFs, which do not contribute to the cross-entropy loss, are minimized, thereby enhancing the model’s robustness against potential backdoor attacks. As a result, Figure 1 (c) and (d) shows that RVPT effectively increases PR by reducing class-irrelevant features, hence successfully defending against state-of-the-art multimodal backdoor attack, even with one shot per class. Moreover, extensive experiments demonstrate our method’s effectiveness across multiple datasets and backdoor attacks.
2 Related Work
Backdoor attacks and defenses on supervised learning. Backdoor attacks have become an increasingly serious security issue because more and more practitioners choose to adopt third-party datasets, platforms, or even backbones to reduce costs. Research on backdoor attacks has primarily focused on designing triggers that enhance stealthiness [6, 46] and effectiveness [31, 36]. For backdoor defenses, researchers employ various strategies to mitigate backdoor attacks, including: (1) pre-processing defense [45] (2) pre-training defense [5] (3) post-training defense [63, 51] (4) test-time defense [14]. Specifically, our work aligns with the post-training defense category, which assumes that defenders do not have access to the poisoned training data and must eliminate the backdoor threat that has already been implanted into the model using only a limited number of clean samples. There are two main strategies for post-training defense. (1) Pruning-based Methods [30, 52]: These approaches focus on identifying the most suspicious neurons and removing them. (2) Fine-tuning-based Methods [49, 63, 35, 26]: These methods purify the compromised model by directly tuning it with clean data through various algorithms.
Backdoor attacks and defenses on CLIP. Although vision-language models (VLMs) [41, 20, 29] have demonstrated significant improvements across various fields, recent research [3, 58] has revealed vulnerabilities of contrastively pre-trained VLMs to backdoor attacks, drawing substantial attention in the field of backdoor attacks and defenses on CLIP-like models. In terms of attacks, BadCLIP [28] optimizes visual trigger patterns within a dual-embedding guided framework, rendering the attack undetectable. Moreover, many backdoor attack methods on multimodal large language models [32, 33, 27, 38, 34] have been proposed recently. Defense strategies can be primarily categorized into three temporal categories: pre-training defense and post-training defense. For pre-training defense [57, 19, 56], SAFECLIP [57] prevents learning from poisoned examples by dynamically dividing the training data into the safe and risky sets by cosine similarity and only applying uni-modal contrastive loss to different modalities of the risky set. For fine-tuning defense [2, 55, 22], CleanCLIP [2] reduces the impact of spurious relationships of backdoor attacks by enforcing the model to realign representations for independent modalities.
Prompt learning for vision-language models.  Lately, various fine-tuning methods have been introduced to further enhance the impressive capabilities of vision-language models (VLMs) across diverse downstream tasks [23, 18, 62, 54], with prompt learning [61, 42, 59] emerging as a particularly efficient and effective approach. Specifically, in the area of natural language processing, prompt learning regards the text prompt, e.g., “a photo of a <CLS>.” as continuous learnable parameters and optimizes it using only a few examples from the downstream dataset. CoOp [61] first introduces prompt learning to VLMs and proposes a method that optimizes the context of the text prompt while keeping the class name tokens fixed. VPT [21] adjusts the visual features by concatenating the image or image features with the learnable prompt. Besides, BadCLIP [1] injects backdoor during the prompt learning stage, enabling the malicious alteration of text features through trigger-aware prompts for a powerful attack. Despite the popularity and versatility of prompt learning for enhancing the model’s robustness [22, 24, 7, 60, 10], limited research has focused on its application in defending against multimodal backdoor attacks. In particular, current post-training backdoor defense methods for VLM mainly focus on tuning model parameters to mitigate spurious relationships, neglecting the potential of prompt learning. Therefore, We propose Repulsive Visual Prompt Tuning, aiming to fill the research gap by incorporating learnable parameters into the features to defend against multimodal backdoor attacks while freezing the model’s parameters.
3 The Proposed Approach
Figure 2: Illustration of RVPT. RVPT concatenates features of later layers with tunable visual prompts while keeping the CLIP’s pre-trained parameters frozen. To optimize the visual prompts, RVPT employs both cross-entropy (CE) loss and feature-repelling (FR) loss. The FR loss minimizes the mean cosine similarity between the prompted features and the original features, effectively filtering out class-irrelevant features that do not contribute to the CE loss. Meanwhile, the CE loss ensures the clean accuracy of the prompted model.
In this section, we first outline the basic settings of the threat model and defender’s goal in Part 3.1, followed by a detailed illustration of how the adversary launches backdoor attacks on CLIP in Part 3.2. Eventually, in Part 3.3, we present Repulsive Visual Prompt Tuning, designed to defend backdoored CLIP against multiple multimodal backdoor attacks.
3.1 Basic Settings
Threat model.  We assume the adversary has access to the training dataset and can produce a backdoored model by injecting carefully crafted poisoned data into the original clean dataset. For multimodal attacks, the poisoned data usually consists of images with the trigger pattern and the proxy captions for the target class. Once trained on the poisoned dataset, the model misclassifies inputs containing the specific trigger pattern into the predefined target class while maintaining correct classification for all other benign inputs. Afterward, the adversary may release the backdoored model online, raising significant security concerns.
Defender’s goal.  We suppose that the defender has access to the backdoored model and a limited set of clean downstream data. The defender’s goal is to obtain a new model utilizing both the backdoored model and the available clean data, wherein the backdoor effects are mitigated while preserving the benign performance of the new model on the downstream clean data. Furthermore, the defender has full control over the recovering process of the backdoored model.
3.2 Backdoor Attacks on CLIP
In general, the CLIP model comprises a visual encoder $\mathcal{V}(\cdot)$, a textual encoder $\mathcal{T}(\cdot)$, and a projection matrix $\boldsymbol{P}$ that projects both the visual and textual representations into the joint feature space. The training dataset contains image-text pairs represented as $\mathcal{D}=\left\{\left(\boldsymbol{x}_{i},\boldsymbol{t}_{i}\right)\right\}_{i=1}^{N}$, where $\boldsymbol{t_{i}}$ serves as a short caption for the corresponding image $\boldsymbol{x_{i}}$.
During the attack on CLIP in the training stage, adversaries typically inject a small portion of poisoned image-text pairs denoted as $\widetilde{\mathcal{D}_{p}}=\{(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{i})\}_{i=1}^{M}$, with $\widetilde{\boldsymbol{x}}_{i}=\boldsymbol{x_{i}}+\Theta$. $\widetilde{\boldsymbol{x}}_{i}$ represents a poisoned image obtained by modifying the original image $\boldsymbol{x}_{i}$ with the trigger pattern $\Theta$, while $\widetilde{t}_{i}$ denotes the proxy caption for the target class $y_{t}$. Consequently, the original benign training dataset $\mathcal{D}$ can be poisoned, resulting in $\widetilde{\mathcal{D}}=\{\widetilde{\mathcal{D}_{p}}\cup\mathcal{D}\}$. During training, CLIP employs contrastive learning to maximize the cosine similarity of positive pairs, expressed as $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{i})=\cos(\boldsymbol{P}\mathcal{V}(\widetilde{\boldsymbol{x}}_{i}),\boldsymbol{P}\mathcal{T}(\widetilde{\boldsymbol{t}}_{i}))$, and minimize the similarity of negative pairs $\phi(\widetilde{\boldsymbol{x}}_{i},\widetilde{\boldsymbol{t}}_{j\neq i})$. To clarify, the selected training image-text pairs $(\widetilde{\boldsymbol{x}},\widetilde{\boldsymbol{t}})$ belong to $\widetilde{\mathcal{D}}$.
Upon training with $\widetilde{\mathcal{D}}$, the encoders of CLIP will be backdoored, denoted as $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$. For the backdoored version of CLIP, the trigger pattern $\Theta$ will have a spurious correlation with the proxy caption of the target class $y_{t}$. Thus, during the inference stage, when the model encounters the attacked images containing the trigger pattern, the predicted posterior probability of the image being classified as the target class $y_{t}$ will increase significantly. In the meantime, the backdoored CLIP performs normally on benign inputs.
3.3 Repulsive Visual Prompt Tuning
In this part, we present Repulsive Visual Prompt Tuning (RVPT), a streamlined yet powerful framework designed to defend against multimodal backdoor attacks. Our key insight is to reduce trigger features of the attacked images by restraining the encoding of class-irrelevant features at each layer of the CLIP visual encoder. This approach will increase CLIP’s perturbation resistivity, which is positively correlated with CLIP’s backdoor robustness.
Formally, assume the defender is given the backdoored CLIP, denoted by $\{\widetilde{\mathcal{V}}(\cdot),\widetilde{\mathcal{T}}(\cdot)\}$ alongside a small set of clean labeled samples $\{\boldsymbol{x}_{i},y_{i}\}^{N}_{i=1}$, $y_{i}\in\mathcal{Y}=\{1,\dots,C\}$. In particular, our focus is on CLIP, whose visual encoder is Vision Transformer [11]. For the clean input $\boldsymbol{x}_{i}$, the feature representation at the $d$-th layer of CLIP’s visual encoder is defined as $\boldsymbol{f}^{d}_{i}=[\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i}]$, where $\boldsymbol{c}^{d}_{i}\in\mathbb{R}^{1\times d_{v}}$ denotes the class embedding that will be projected to the vision-language joint space and $\boldsymbol{e}^{d}_{i}\in\mathbb{R}^{M\times d_{v}}$ are the patch embeddings.
Assume the visual encoder has $L$ layers, then starting from the $D$-th layer, we introduce $L-D$ repulsive visual prompts, each of length $b$, defined as $\{\boldsymbol{p}^{k}\in\mathbb{R}^{b\times d_{v}}\}_{d=D}^{L-1}$. These prompts are concatenated with the features from each subsequent block recursively, following the formulation:
$[\boldsymbol{c}^{d+1}_{i},\boldsymbol{e}^{d+1}_{i},\_]=\mathcal{V}^{d}([\boldsymbol{c}^{d}_{i},\boldsymbol{e}^{d}_{i},\boldsymbol{p}^{d}_{i}]),d\!=\!D,\dots,L\!-\!1.$ (2)
We then regard $\boldsymbol{c}^{d}_{i},d\in[D+1,L]$ as the feature to repel and apply feature-repelling (FR) Loss to minimize the cosine similarity between it and its counterpart $\boldsymbol{\sigma}^{d}_{i}$ in the original fixed backdoored visual encoder. For a batch of $N_{b}$ examples $\{\boldsymbol{x}_{i},y_{i}\}^{N_{b}}_{i=1}$, the FR loss is computed as:
$\mathcal{L}_{\mathrm{FR}}=\frac{1}{L-D}\sum_{i=1}^{N_{b}}\nolimits\sum_{d=D+1}^{L}\nolimits\cos(\boldsymbol{c}^{d}_{i},\boldsymbol{\sigma}^{d}_{i}).$ (3)
Additionally, we also optimize the repulsive prompt with the cross-entropy (CE) loss to avoid repelling the class-relevant features and maintain clean accuracy, formulated as:
$\mathcal{L}_{\mathrm{CE}}=-\sum_{i=1}^{N_{b}}\nolimits\mathrm{log}(\frac{\widetilde{\phi}(\boldsymbol{x}_{i},T(y_{i}))/\tau}{\sum_{c=1}^{C}\widetilde{\phi}(\boldsymbol{x}_{i},T(c))/\tau}),$ (4)
where $\tau$ denotes a temperature parameter to control the logit distribution, $T(\cdot)$ is the proxy caption of the input class, e.g., “a photo of a <CLS>.” and $\widetilde{\phi}(\boldsymbol{x}_{i},\boldsymbol{t}_{j})=\cos(\boldsymbol{P}\widetilde{\mathcal{V}}(\boldsymbol{x}_{i}),\boldsymbol{P}\widetilde{\mathcal{T}}(\boldsymbol{t}_{j}))$ is the function of cosine similarity computed with outputs by the recovering encoders.
Thus, the overall loss for RVPT is defined as:
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\alpha\mathcal{L}_{\mathrm{FR}},$ (5)
where $\alpha$ is the balancing factor controlling FR loss strength.
For further explanation, RVPT can be interpreted through the lens of natural selection [8]. To be specific, FR loss creates a challenging environment for visual prompt tuning to optimize CE loss. In order to simultaneously minimize CE loss and maximize the discrepancy between the prompted features and original features, the visual prompts must learn to select the most suitable class-relevant features while discarding those that are class-irrelevant. Ultimately, this process leads to a more compact visual encoding mechanism.
4 Experiments
[…]
Method BadNet Blended Label Consistent ISSBA WaNet TrojVQA BadCLIP
ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA
No defense 87.49 56.34 99.22 56.08 86.14 55.97 70.19 55.69 87.58 53.05 99.72 56.62 99.83 60.63
CleanCLIP 56.55 52.77 34.33 54.22 62.97 52.74 51.85 54.06 70.03 52.37 79.69 54.90 67.53 53.42
Linear Probe 5.29 56.83 1.78 57.12 1.94 56.97 0.09 56.79 1.31 55.75 - - 99.70 59.33
RVPT 1.84 58.09 0.02 57.97 1.14 57.92 0.01 58.18 0.09 56.22 3.33 58.12 2.76 61.72
[…]
Method Caltech101 OxfordPets
BadNet Blended WaNet BadNet Blended WaNet
ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA
No defense 88.85 91.07 91.28 90.69 80.91 88.63 93.89 80.36 100 71.24 99.91 75.81
CleanCLIP 29.40 88.26 15.72 90.25 24.66 91.10 72.57 70.45 32.98 73.25 88.06 75.11
Linear Probe 1.47 92.53 0.53 93.37 1.03 91.74 9.02 75.75 3.71 70.43 3.70 74.89
RVPT 0.04 93.55 0 93.40 0.28 93.33 0.64 86.02 2.97 83.62 2.68 85.49
4.1 Experiment Settings
Models and benchmarks. Following the settings of [2], to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M [41] using 500K image-text pairs from the CC3M dataset [44], out of 1500 are poisoned with various types of backdoor attacks: BadNet [15], Blended [6], Label Consistent [46], ISSBA [25], WaNet [37], TrojVQA [48] and BadCLIP [28]. For BadNet, Blended, and WaNet, the target class is banana, accordion, or samoyed. For Label Consistent, ISSBA, WaNet, TrojVQA and BadCLIP, the target class is banana. Our defense mainly utilizes ImageNet [9] to defend against backdoor attacks whose target class is banana, Caltech101 [13] for accordion and OxfordPet [39] for samoyed. We compare RVPT with Linear Probe and CleanCLIP [2]. The implementation details of the aforementioned backdoor attack methods, backdoor defense methods, and benchmarks will be thoroughly illustrated in the supplemental material. Additionally, we report the zero-shot results of CleanCLIP and RVPT in the following experiment. The results of combinations of defense methods, such as CleanCLIP + Linear Probe, RVPT + Linear Probe, CleanCLIP + RVPT will also be provided in the supplemental material.
Evaluation criteria. For backdoor attacks targeting the class of banana, accordion, or samoyed, we evaluate the defense methods utilizing the test sets of ImageNet, Caltech101, and OxfordPets separately. The evaluation metrics include Clean Accuracy (CA) and Attack Success Rate (ASR). CA is the model’s accuracy on clean images, while ASR measures the proportion of poisoned images predicted as the target label. A lower ASR and decent CA stand for a successful defense.
Implementation details. The implementation of our defense method is based on Pytorch [40]. Unless otherwise specified, we will use ViT/B32 [11] as CLIP’s visual encoder. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, we set the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP [41]: (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule.
4.2 The Effectiveness of RVPT
In Table 1 and 2, we compare RVPT with CleanCLIP and Linear Probe against seven attacks on ImageNet and three attacks on Caltech101 and OxfordPets.
Firstly, we can see all attacks achieve a high ASR while maintaining decent CA with no defense, indicating all attacks have been successfully launched.
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance on downstream tasks and requires more computing resources (it needs full parameter tuning and much more image-text pairs). Moreover, Linear Probe performs well against the first five attacks without losing clean accuracy. However, it actually does not remove the backdoor implanted in the visual encoder and performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since CLIP tuned with Linear Probe will only retain visual encoder, we opted not to apply it to defend against TrojVQA, which focused on the multimodal task of visual question answering.
Finally, RVPT successfully defends against all types of backdoor attacks, with a notable performance gain: CA increased by 1% to 2% generally. In particular, it performs better against global trigger patterns, such as Blended, ISSBA, or WaNet. We conjecture the reason behind this phenomenon lies in the fact that the global trigger pattern rarely exists in the downstream dataset, which is easier to be recognized as the class-irrelevant feature and removed by RVPT.
[…]
Target class RVPT
BadNet (woTC) BadNet Blended (woTC) Blended WaNet (woTC) WaNet
banana 2.18 (58.15) 1.84 (58.09) 0.01 (57.73) 0.02 (57.97) 0.09 (55.34) 0.09 (56.22)
samoyed 0.76 (84.25) 0.64 (86.02) 0.03 (81.64) 2.97 (83.62) 1.03 (83.76) 2.68 (85.39)
accordion 2.29 (93.41) 0.04 (93.55) 0 (92.83) 0 (93.40) 0 (92.94) 0.28 (93.33)
[…]
Method accordion samoyed
BadNet Blended WaNet BadNet Blended WaNet
CleanCLIP 29.40 (88.26) 15.72 (90.25) 24.66 (91.10) 72.57 (70.45) 32.98 (73.25) 88.06 (75.11)
RVPT* 5.59 (87.47) 1.02 (88.03) 2.49(89.02) 5.97 (76.90) 0.64 (76.53) 3.42 (76.79)
4.3 RVPT’s Defensive Ability Can Generalize
We evaluate RVPT’s generalization ability in defending backdoor attacks across three scenarios:
  • The target class of the attacks is not in the tuning dataset.
  • The attacks are launched for downstream datasets different from the tuning dataset.
  • The attacks are launched for the cross-domain datasets.
[…]
Method ImageNet-V2
BadNet Blended BadCLIP
No defense 89.06 (49.33) 99.53 (49.04) 99.89 (53.49)
CleanCLIP 62.92 (45.58) 40.57 (47.24) 73.07 (46.20)
RVPT 2.64 (49.35) 0.02 (49.51) 3.43 (52.53)
Method ImageNet-A
BadNet Blended BadCLIP
No defense 92.96 (23.31) 99.92 (24.10) 99.97 (30.80)
CleanCLIP 79.19 (19.25) 71.03 (20.79) 92.27 (20.58)
RVPT 7.67 (14.25) 0.09 (14.69) 12.84 (16.84)
Method ImageNet-R
BadNet Blended BadCLIP
No defense 89.59 (61.24) 98.61 (61.97) 99.69 (65.49)
CleanCLIP 68.66 (58.79) 44.31 (60.29) 72.32 (60.29)
RVPT 3.56 (53.08) 0.08 (53.02) 6.75 (57.37)
Method ImageNet-S
BadNet Blended BadCLIP
No defense 93.81 (36.43) 96.48 (36.88) 99.88 (40.19)
CleanCLIP 76.37 (33.38) 14.26 (34.33) 69.80 (34.46)
RVPT 1.92 (31.16) 0 (30.76) 1.67 (34.06)
Performance on the emerging target class.  We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its open-vocabulary defense capability for unseen target classes, and the results are shown in Table 3. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 3 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully identified and removed the distinctive pattern. For BadNet, which is harder to resist, ASR increases slightly, though still within a reasonable range. Moreover, we can see that CA maintains other than OxfordPets, which decreases by around 2%. We believe the reason is that this fine-grained dataset only contains 37 classes. The absence of 3% of the data can significantly impact accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
Performance on cross-dataset transfer.  To evaluate RVPT’s generalization of defensive ability across datasets, we fine-tuned RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 4. For all three attacks, although ASR increases slightly, it remains at a low level. Additionally, RVPT’s accuracy decreases occasionally. We conjecture the reason is that ImageNet does not share some of the class-irrelevant features (CIFs) with Caltech101 or OxfordPets. In some cases, The CIFs in ImageNet may be the discriminative features in them. Nevertheless, the CIFs of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset is one reason for the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still defends better than CleanCLIP.
Performance on cross-domain transfer.  To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 [43], ImageNet-A [17], ImageNet-R [16], and ImageNet-S [50]. The result is shown in Table 5. These three attacks can generalize well on the four benchmarks. We observe that RVPT will experience a loss of recognition and defense performance with domain shifts, especially in ImageNet-A. We thus conjecture that the CIFs of cross-domain images vary more than real-world images. Notably, even with cross-domain images, RVPT still outperforms CleanCLIP and significantly decreases ASR.
4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT.
Number of shots per class.  How many shots are needed to guarantee a successful defense? From Figure 3 (a), we surprisingly find that only 1 shot per class can make the ASR of Blended less than 0.02% and significantly decrease the ASR for other attacks. Moreover, increasing the number of shots further enhances the defense. For BadNet and BadCLIP, 16 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving, which might be caused by the limited repelling capabilities of the visual prompts.
Balancing factor $\alpha$.  We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3 (b), we observe that ASR sharply decreases while CA sharply decreases as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of class-relevant features (the environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small balancing factor, like 2, can achieve very low ASR while preserving relatively strong CA for every evaluated dataset.
Visual context length.  We show RVPT’s performance in Figure 3 (c) with changing visual context length $b$. It indicates that both ASR and CA are optimal at a context length of approximately 25 to 50 tokens. Beyond this optimal range, CA decreases due to overfitting caused by the increasing number of parameters. Moreover, with increasing length, the defensive ability is bounded.
Depth of the visual prompt.  Figure 3 (d) illustrates the performance as the depth of the features into which the visual prompts are embedded is varied. It is important to note that “depth” here refers to the number of layers counted from the model’s output feature, as the later layers are more significant for defending against backdoor attacks. This phenomenon can be observed in Figure 3 (e), which reflects the influence of depth measured from the input features. The rationale behind this observation lies in the nature of class-irrelevant features encoded in the earlier layers, which tend to be sparse and intermixed with desirable features. Consequently, repelling these class-irrelevant features at an earlier stage is difficult and may inadvertently compromise the encoding of discriminative features. Moreover, the results in Figure 3 (d) indicate the improving performance as the number of the prompted layers increases. Particularly, when the depth reaches approximately six layers, both CA and ASR attain their optimal values and will not deteriorate.
Margin of the repelling strength.  To evaluate the potential risks of over-repelling, we set a margin $m$ of FR Loss $\mathcal{L}_{\mathrm{FRM}_{i}}$ for each sample $\boldsymbol{x}_{i}$:
$\mathcal{L}_{\mathrm{FRM}_{i}}=\mathrm{max}(\mathcal{L}_{\mathrm{FR}_{i}},m).$ (6)
A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion. The performance of RVPT of different $m$ is shown in Figure 3 (f). We observe that both ASR and CA are sensitive with $m$ and will increase with larger $m$. In order to achieve superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
FR loss.  To assess the backdoor robustness introduced by FR loss in deep visual prompt tuning, we compare the ASR of RVPT and VPT against BadNet attack and Blended attack in different scenarios, as shown in Figure 4. The results can also be regarded as an ablation analysis of RVPT, which indicates that while VPT alone reduces ASR to some extent, FR loss significantly enhances the defensive capability against backdoor attacks across various conditions.
[…]
4.5 Qualitative Analysis
Figure 5: Attention map for (a) original image and (b) poisoned image. The second column shows the attention map of the backdoored visual encoder attacked by BadCLIP. The third and fourth columns show the attention map of RVPT’s patch embeddings and the prompt embeddings. RVPT effectively reduces the over-attention of both the class-irrelevant features and the trigger pattern.
Attention map.  We plot the attention maps for the backdoored encoder and the recovered encoder by RVPT in Figure 5 to observe the areas that the model focuses on. Firstly, for the clean image, RVPT enables the encoder to focus more on the useful features of the bird. In contrast, the backdoored encoder also attends to irrelevant features that humans cannot interpret. Secondly, for the poisoned image, although RVPT has noticed the trigger pattern, unlike the backdoored encoder, it will not draw all the attention to it.
Figure 6: The t-SNE [47] plots for the representations of clean (dotted) and poisoned (star-shaped) images. The model is attacked by BadCLIP. We can see that RVPT successfully breaks the clustering effect of the poisoned image features in the embedding space. This clustering effect of poisoned representations was revealed in [2].
t-SNE.  We draw the t-SNE plots [47] in Figure 6 to observe the distribution of the image representations of different visual encoders. The plots show that image representations encoded by the backdoored CLIP and VPT-defended CLIP are clustered, indicating that these models focus solely on the trigger pattern, ignoring the information from the original image. On the contrary, RVPT not only disrupts this cluster but also restores the poisoned image representations, bringing them closer to their original counterparts. This observation suggests that after RVPT, the trigger pattern will have a minimal effect on the visual representations.
5 Conclusion
In this paper, we, for the first time, investigated how to utilize limited clean samples and tunable parameters to defend a backdoored multimodal contrastively pre-trained model (i.e., CLIP). We empirically showed that the excessive encoding of class-irrelevant features causes CLIP’s poor resistivity to input perturbations, resulting in its vulnerability to backdoor attacks. To reduce the class-irrelevant features, we proposed a novel method called Repulsive Visual Prompt Tuning. This method leverages our designed feature-repelling loss to minimize the mean cosine similarity between the prompted features and the original features while optimizing cross-entropy loss to maintain clean accuracy. Comprehensive experimental results on multiple attacks and benchmark datasets demonstrated the effectiveness of our proposed method.
Supplemental Material
Organization of the Supplemental Material
  • Section A provides more results of different perturbations and shows the negative correlation between PR and ASR.
  • Section B provides more detail in the experiment’s settings of backdoor defense and attack methods.
  • Section C shows the experiment results of combining the proposed method and other backdoor defense methods.
  • Section D contains more experiments of ablation study, such as the distance measure function, the permutation of the repulsive prompts, etc. It also contains the setting of experiments that ablates the FR loss.
  • Section E compares computational expenses between our method and CleanCLIP.
A More Analysis of Perturbation Resistivity
A.1 More Types of Perturbation
We evaluate the perturbation resistivity of four models: ViT/B32 of OpenAI CLIP, ViT supervised exclusively on ImageNet, RVPT with 1 shot on ImageNet, and RVPT with 16 shots on ImageNet against various input perturbations on the ImageNet validation set. The visual illustration is shown in Figure 1, and the results of the PR values for different models are in Table 1. The results in Table 1 show that CLIP has generally much less PR value against various perturbations than traditionally supervised ViT, and RVPT can consistently increase its PR value across all scenarios. Even with 1 shot, RVPT can effectively increase CLIP’s PR value and thus successfully defend it from various backdoor attacks.
Figure 1: Examples of the experimented perturbation. We adopt various perturbations to verify CLIP’s intrinsically low PR value.
Table 1: Perturbation resistivity of the four models across various perturbations. We can see the PR value of CLIP ViT is less than supervised ViT, while RVPT significantly improves its PR.
Perturbation ImageNet
Supervised ViT CLIP ViT VPT - 16 shot RVPT - 1 shot RVPT - 16 shot
Gaussian noise (std=0.01) 0.959 0.572 0.737 0.942 0.965
Gaussian noise (std=0.03) 0.951 0.370 0.583 0.906 0.931
Gaussian blur (kernel size=3) 0.976 0.882 0.886 0.986 0.992
Gaussian blur (kernel size=5) 0.975 0.819 0.890 0.977 0.986
Color jitter 0.946 0.832 0.887 0.974 0.983
Horizontal flipping 0.971 0.950 0.968 0.991 0.994
BadNet attack 0.976 0.861 0.954 0.982 0.991
Blended attack 0.975 0.855 0.915 0.980 0.990
BadCLIP attack (trigger pattern) 0.975 0.871 0.550 0.974 0.988
A.2 The relationship between PR and ASR
We evaluate the relationship using two kinds of perturbations, including Gaussian noise and non-adversarial trigger patterns. From Figure 2, we can see a clear, strong negative correlation between PR and ASR for different types of backdoored models.
Figure 2: The correlation between PR and ASR. By varying the prompted depth in RVPT, we recorded the PR and ASR for different depths. The result shows a clear negative correlation between PR and ASR. In (a), we use Gaussian noise whose standard deviation is 0.01; In (b), we use Blended to perturb the images fed into the CLIP infected by BadNet. For CLIP infected by BadCLIP and Blended, we use the perturbation of Random. Therefore, it justifies the motivation to reduce PR.
B Detailed Settings
B.1 Detailed Settings of Backdoor Attacks
In our experiments, we evaluate seven prominent backdoor attack methods: BadNet [15], Blended [6], BadCLIP [28], ISSBA [25], Label Consistent [46], WaNet [37], and TrojVQA [48]. Below, we provide a detailed description of each method:
  • BadNet, a foundational approach to backdoor attacks in deep learning, generates poisoned samples by embedding a small, randomly placed patch into images and modifying their labels to match the target class. In our implementation, the patch size is set to 16 pixels.
  • Blended enhances the stealth of backdoor attacks by focusing on trigger design. This method linearly blends the trigger with the original image, creating an almost imperceptible backdoor. We adopt a blending ratio of 0.2 for the trigger to ensure minimal visibility to the human.
  • BadCLIP targets CLIP models with a backdoor attack that optimizes visual trigger patterns using a dual-embedding guided framework. This method ensures that the attack remains undetectable by leveraging both text and image embeddings. We follow the parameter configurations in the original paper to implement BadCLIP.
  • ISSBA introduces a highly covert attack by generating sample-specific triggers. These triggers are embedded into benign images using an encoder-decoder architecture, encoding a predefined string into the images. Moreover, the ciphertext is the string ‘Stega!!’.
  • Label Consistent improves backdoor attack stealth from the perspective of label manipulation. It employs generative models or adversarial perturbations to selectively poison images that belong to the target class, maintaining consistency between poisoned images and their original class labels to escape human inspection. We use BadNet-LC in our experiment.
  • WaNet proposes warping-based triggers to make the attack less noticeable to humans. Specifically, we use control grid size k = 224 and warping strength s = 1 and train models without the noise mode.
  • TrojVQA proposes a dual-key backdoor attack that sets triggers on both modalities. To evaluate RVPT under backdoored generative VLMs, we use TrojVQA as the backdoor attack and use the following settings. Firstly, we optimize the visual triggers for the CLIP visual backbone with the description “This is a yellow banana.”. Specifically, We set the patch size to 16$\times$16 and locate the patch in the middle of the image. Then, we randomly corrupt the text for poisoned samples into sentences that describe the target class of banana. Moreover, we add the trigger word “SUDO ” in front of each sentence to make sure two keys for one backdoor. In the inference stage, the optimized trigger is added to the image, and “SUDO ” is added to the beginning of the prompt.
We employ the AdamW optimizer with an initial learning rate of 1e-5 for all the backdoor attack methods except 1.5e-5 for WaNet. The learning rate follows a cosine decay schedule over five epochs, and we set the batch size to 128. The visualization of each trigger pattern is in Figure 3.
Figure 3: Visualization of the triggers of various experimented attacks.
B.2 Detailed Settings of Backdoor Defenses
CleanCLIP  CleanCLIP [2] provides a defense against backdoor attacks in multimodal contrastive learning by optimizing the integration of multimodal contrastive and unimodal self-supervised losses, relying on a limited amount of clean data. It is important to note that CleanCLIP utilizes ResNet-50 as the backbone of its visual encoder. In contrast, our work adopts the Vision Transformer (ViT-B/32) as the visual encoder, and we have accordingly adjusted the parameters to fit this architecture. Specifically, we randomly selected 100,000 image-text pairs from the CC3M dataset for fine-tuning. The learning rates were set to 5e-6 for BadNet, TrojVQA, Blended, Label Consistent, and BadCLIP, while for WaNet and ISSBA on ImageNet-1K, they were set to 3e-6. We used a batch size of 64 and fine-tuned the models over 10 epochs. Notably, we did not focus solely on reducing the attack success rates by manipulating learning rates; instead, we ensured that the clean accuracy of the fine-tuned model was consistently maintained throughout.
Linear Probe  Linear Probe [3] adds a tunable linear layer on top of the visual encoder, which seems efficient and effective in defending backdoor attacks. It does not truly purify the hidden trigger features. Consequently, it becomes ineffective against state-of-the-art multimodal attacks, whose trigger pattern aligns with the local image features [28]. This paper uses the same settings in [11] to adopt ViT-B/32 using the linear layer. Specifically, we use SGD as the optimizer and use the constant learning rate of 0.01 and momentum of 0.9. The batch size is 32, and the total training epoch is 40.
B.3 Detailed Settings of Datasets
This paper assesses ASR and CA using three downstream datasets: ImageNet1K [9], Caltech101 [13], and OxfordPets [39]. For cross-domain evaluation, ImageNet-V2 [43], ImageNet-A [17], ImageNet-R [16], ImageNet-Sketch [50] are also evaluated. Additionally, CleanCLIP selects clean image-text pairs from CC3M [44] for fine-tuning the backdoored CLIP model. Below are detailed descriptions of these datasets:
  • ImageNet1K includes 1,000 classes with over a million images, presenting a complex, large-scale dataset for image classification challenges.
  • Caltech101 comprises 101 object classes and one background category, with each class containing between 40 and 800 images. It is widely used for evaluating models on fine-grained classification and image recognition.
  • OxfordPets is a fine-grained dataset with 37 categories and approximately 200 images per class. The images display wide variations in scale, pose, and lighting.
  • ImageNet-V2 was developed to assess model generalization under temporal shifts. It closely mirrors the original ImageNet’s data collection and annotation processes, using new samples sourced from the Internet.
  • ImageNet-A is designed to compile images that present notable challenges to deep learning models, often causing misclassification and thereby highlighting potential model vulnerabilities.
  • ImageNet-R features “non-real” imagery, such as artworks, cartoons, and graphical representations, which differ significantly in visual style and form from the natural images found in the original ImageNet.
  • ImageNet-Sketch contains hand-drawn sketches of the same categories as ImageNet, characterized by a focus on lines and shapes rather than colors and textures, presenting a distinct departure from photographic imagery.
  • CC3M, or Conceptual Captions, includes around 3.3 million images paired with captions. Unlike other datasets with curated annotations, CC3M’s image descriptions are sourced from web Alt-text attributes, thus offering a broader range of descriptive styles and contexts.
C Results of Defense Combination
Table 2 shows that RVPT still shows superiority when defending against multiple backdoor attacks on ImageNet. Moreover, when RVPT is combined with a Linear Probe, the model overfits the dataset with significantly dropping accuracy. In summary, CleanCLIP + RVPT is the best choice to defend against backdoor attacks.
[…]
             Method              ImageNet
             BadNet              Blended              BadCLIP
             CP + LP              0.60 (54.64)              0.24 (56.24)              53.30 (55.79)
             RVPT + LP              0.26 (51.39)              0.04 (50.26)              0.09 (49.53)
             CP + RVPT              0.54 (55.65)              0.03 (57.00)              0.37 (56.45)
D More Experiments for Ablation Study
D.1 Ablation Study of the Distance Measure Function
Instead of using cosine similarity, we experimented with maximizing the L1 and L2 distances of the normalized features, as shown in Table 3. However, both L1 and L2 losses demonstrate suboptimal performance.
[…]
           Method            ImageNet
           BadNet            blended            BadCLIP
           L1 distance            46.50 (60.20)            16.84 (60.14)            0.04 (53.72)
           L2 distance            48.72 (60.40)            77.44 (60.30)            94.73 (63.65)
           cos similarity            1.84 (58.09)            0.02 (57.97)            2.76 (61.72)
D.2 Ablation Study of the Permutation of the prompts and FR loss
To evaluate the best permutation of the prompts and FR loss, we fix the number of repulsive prompts and repelled features and design 4 kinds of permutations: Split, Overlap, Intertwine, Mix in Figure 4 and show the results in Table 4. From the result, we can conclude that repulsive prompts and repelled features should be close to the output to obtain optimal performance of both recognition and defense.
Figure 4: Visualization of the different permutations: red bars indicate that the layer’s feature is prompted, blue bars indicate that the layer’s feature is repelled, and purple bars indicate that the layer’s feature is both prompted and repelled.
[…]
               Permutation                ImageNet
               BadNet                blended                BadCLIP
               Split                7.89 (53.79)                0.03 (53.60)                12.49 (57.99)
               Overlap                1.38 (57.20)                0.05 (57.20)                7.16 (61.24)
               Intertwine                14.72 (55.47)                0.02 (55.44)                27.02 (59.98)
               Mix                9.17 (57.33)                0.02 (56.92)                9.48 (61.02)
D.3 Ablation Study of the Handcrafted Prompt
We experiment with the handcrafted prompt of “<CLS>” and “###### <CLS>” and show the result in Table 5. From the result, we can conclude that RVPT still successfully defends against backdoor attacks with different handcrafted prompts.
[…]
           Handcrafted Prompt            ImageNet
           BadNet            blended            BadCLIP
           “<CLS>”            2.10 (58.05)            0.02 (58.00)            3.55 (61.73)
           “###### <CLS>”            1.91 (57.62)            0.01 (57.47)            3.76 (61.36)
           “a photo of a <CLS>”            1.84 (58.09)            0.02 (57.97)            2.76 (61.72)
D.4 Ablation Study of the Visual Encoder Architecture
To evaluate RVPT across different architectures, We poisoned different architectures of CLIP using the same hyper-parameters and show the result in Table 6. In particular, for the sake of fairness, we keep the RVPT training setting for ViT-B/16 and ViT-L/14 the same as for ViT-B/32. The result shows that all attacks have been launched successfully, so we can conclude that RVPT still successfully defends against backdoor attacks in various visual encoder architectures.
[…]
            ImageNet
            Method             BadNet             Blended             BadCLIP
            ViT-B/16
            No defense             99.61 (68.04)             98.30 (67.84)             99.67 (68.09)
            Linear Probe             7.11 (67.52)             0.21 (67.30)             97.73 (67.60)
            RVPT             0.66 (68.51)             0.02 (68.32)             3.38 (68.42)
            ViT-L/14
            No defense             98.76 (73.88)             99.02 (74.25)             99.87 (74.68)
            Linear Probe             23.48 (72.31)             0.65 (72.35)             99.35 (72.35)
            RVPT             2.08 (73.60)             0.05 (73.26)             0.59 (74.04)
D.5 Settings for Experiment of Ablation Study
First of all, we use models backdoored by Blended attack and BadNet attack separately to conduct all the evaluations.
  • For the evaluation of in-dataset ASR, we use the infected model whose target class is banana and tune it with 16-shot training samples on ImageNet;
  • For the evaluation of open-vocabulary ASR, we tune the infected model whose target class is banana with the few-shot training samples in ImageNet, which does not have the images of the target class.
  • To evaluate cross-dataset ASR, we use infected models whose target class is accordion to be tuned on ImageNet and evaluate their ASR on Caltech101.
  • For the evaluation of cross-domain ASR, we evaluate the ASR of ImageNet-tuned models on ImageNet-Sketch.
E Computational Expenses
We conducted the main experiments on a single NVIDIA 3090 using a half-precision data type and recorded one training process in Table 7. The results indicate that our method substantially reduces computational costs, enhancing defense efficiency.
Table 7: Computational expense comparison between RVPT and CleanCLIP.
Method Training time GPU memory used Tunable parameters Training samples
CleanCLIP 4:46:21 13854 MB 126 M 100K
RVPT on ImageNet 45:05 3382 MB 0.34 M 16 K
RVPT on OxfordPets 2:19 1010 MB 0.34 M 0.6 K